# Doping-Driven Modulation of Spin–Orbit Coupling, Spin Textures, and Rashba–Edelstein Response in Chiral Tellurium: A First-Principles Study

### v2

@@ III.3 Spin texture
The spin textures projected onto the $k_{y}$$k_{z}$ plane reveal marked differences between pristine tellurium (Te) and its doped variants (S-, Sb-, and Se-substituted Te). In pristine Te, the spins exhibit a predominantly radial pattern emanating from (or converging toward) high-symmetry points in the Brillouin zone, consistent with the chiral-induced spin textures previously reported for elemental Te. In particular, the spins in pristine Te appear to tilt systematically as one moves along $k_{y}$ or $k_{z}$, a feature that has been attributed to strong spin-orbit coupling (SOC) combined with the crystal’s helical arrangement of atoms. This observation aligns with earlier experimental and theoretical studies of spin textures in chiral Te, which showed similar radial or nearly collinear patterns in momentum space [25, 8]. The net outcome is that, even without an external magnetic field, spin polarization in pristine Te is locked to the crystal momentum in a manner driven by Te’s chirality.
@@
Upon S substitution, the spin texture retains the general radial character but displays a slight enhancement of the spin polarization angles in certain regions of the Fermi surface. In other words, the spins remain roughly outward-pointing, but the degree of out-of-plane tilt varies more strongly across the Brillouin zone. This trend may be attributed to the lighter atomic mass of S relative to Te, subtly modifying the effective SOC strength and shifting the conduction or valence bands. Such doping-induced tuning of the spin texture is in line with the view that substituting lighter elements can reduce the overall SOC, thereby sharpening any radial or tangential spin patterns within specific energy pockets [50]. Notably, the S-doped sample maintains the underlying chirality of the host lattice, so the qualitative features of the spin texture remain intact.
@@
By contrast, the Sb-doped system shows a somewhat more pronounced spin polarization, especially at larger $|k_{y}|$ or $|k_{z}|$. This observation suggests that heavier isoelectronic doping can amplify SOC-driven spin splitting, leading to larger momentum-dependent variations in spin orientation. The increased atomic mass of Sb relative to Te has often been linked to enhanced spin-orbit coupling, which can in turn reinforce the chiral spin-moment locking in the conduction bands [51]. In the present calculations, the Sb-induced band modifications appear primarily near the Fermi level, thus making the spin texture more conspicuously tilted in those momentum-space regions that dominate conduction. The net result is a doping strategy that can, in principle, boost the spin signals relevant for spintronics applications, in keeping with prior theoretical predictions on chiral semiconductors doped with heavier elements [52].
@@
The Se-doped Te sample, finally, shows a spin texture that is intermediate between the S and Sb cases, likely reflecting the atomic mass and electronic configuration of Se. Compared to S-doped Te, the Se-doped compound exhibits slightly stronger SOC, manifested as increased tilting in certain $(k_{y},k_{z})$ sectors. However, it lacks the more dramatic spin-splitting enhancements seen in Sb doping. Such subtle shifts in spin angles and magnitudes may prove critical if one aims to engineer specific transport properties—such as persistent spin helices or large charge-to-spin conversion efficiencies—in a doping-dependent fashion [53]. The combination of these four systems underscores how different dopants systematically tune the interplay between chirality and SOC, allowing one to either reinforce or diminish the radial-type spin texture.
@@
Taken together, these results corroborate the emerging consensus in the literature that modest doping—whether through lighter or heavier substitution—provides a robust means of controlling spin textures in chiral semiconductors [25, 50]. They also reinforce the notion that chiral crystals such as Te enable tunable spin signals even without magnetism, potentially facilitating magnet-free spintronic devices. Consistent with prior first-principles and experimental reports, the radial or nearly collinear spin textures persist across a broad range of dopant elements, demonstrating that the underlying chirality is remarkably resilient. However, the magnitude and tilt of spin polarization can be selectively modified, which paves the way for designing materials with customized spin-transport characteristics. Future work will extend these analyses to higher doping concentrations or to co-doping scenarios, further mapping out the interplay between doping, band-edge positioning, and spin relaxation times in these technologically promising chiral semiconductors.
@@ III.4 The Rashba–Edelstein (REE)
The Rashba–Edelstein (REE) tensor components for pristine Te and three doped variants (Te:S, Te:Se, Te:Sb) exhibit clear modifications in both magnitude and energy dependence, underscoring how substitutional doping can tune spin-orbit–driven effects in chiral semiconductors. In pristine Te, the three independent components $\chi_{xx}$, $\chi_{yy}$, $\chi_{zz}$ show intermediate magnitudes on the order of $10^{9}$ (in the chosen units) and prominent dispersion across the studied energy window from about $-0.20$ eV up to 00 eV. Notably, $\chi_{xx}$ and $\chi_{yy}$ overlap partially in magnitude but differ in sign over most of the range, whereas $\chi_{zz}$ exhibits a distinct curvature. These features are consistent with the previously reported “radial” or collinear spin textures in elemental Te, in which the interplay between chiral crystal symmetry and spin-orbit coupling (SOC) leads to a sizable Edelstein response over a broad valence-band region [49, 25].
@@
Upon substituting Te with sulfur (Te:S), the overall magnitude of the tensor components increases to around $10^{10}$, indicating that lighter-element substitution can shift and enhance the SOC-driven effects near the Fermi energy. In particular, $\chi_{xx}$ becomes quite large and negative over a significant portion of the valence region, while $\chi_{yy}$ and $\chi_{zz}$ also remain negative but follow somewhat shallower dispersions. This enhancement is consistent with the notion that doping can move the chemical potential to regions where band splittings are pronounced, thereby accentuating the collinear spin polarization [36, 54]. At the same time, the lighter S substitution slightly modifies local bonding environments, which may reduce certain symmetry-allowed band degeneracies and hence enlarge the net REE response.
@@
Doping with selenium (Te:Se) likewise shows magnitudes near $10^{10}$ but with a different balance among the $\chi_{xx}$, $\chi_{yy}$, and $\chi_{zz}$ components. Here, $\chi_{yy}$ acquires its largest negative values around $-0.3$ eV, suggesting that Se substitution emphasizes different orbital characters near the Fermi level compared to S doping. This behavior is in line with earlier studies that highlight how chalcogen substitutions in Te can shift Weyl-like band features or break certain rotational symmetries, thus affecting both spin-splitting and Berry-related phenomena [55]. The capacity to increase or redirect the REE components through chalcogen doping corroborates the hypothesis that lighter dopants can markedly reshape the conduction- and valence-band edges, thereby offering routes to optimize spin accumulation in nonmagnetic chiral semiconductors [56].
@@
In contrast, doping with Sb (Te:Sb) reduces the overall scale of the REE tensor to around $10^{7}$. Intriguingly, $\chi_{xx}$ is positive and relatively large compared to $\chi_{zz}$, which is mostly negative throughout the probed energy range. These trends imply that heavier-element substitution, although it enhances intrinsic SOC strength, can simultaneously push the Fermi level away from the band crossings most responsible for large spin-orbit splitting, leading to a net reduction in the Edelstein coefficients [57, 58]. In this sense, Sb doping might be more suitable for achieving moderate spin-charge conversion while preserving good electronic conduction, rather than maximizing the REE. The interplay between doping type (electron vs. hole doping) and element mass therefore emerges as a key strategy for engineering chiral spin transport, aligning with the broader literature on doping-induced spin accumulation in noncentrosymmetric semiconductors [49].
@@
Overall, these four cases illustrate how modest substitutional doping in chiral Te can selectively tune the sign, amplitude, and energy dependence of the REE tensor, revealing a clear pathway for tailoring spin-based functionalities in chiral semiconductors. The results agree well with prior theoretical and experimental observations that emphasize doping as a pivotal parameter for controlling current-induced magnetization and spin-lifetime effects in topological or near-topological band structures [55, 56]. By situating the Fermi level at different spin-split band extrema—and by exploiting the modified SOC landscapes—one can either amplify spin accumulation (as in S or Se doping) or preserve moderate REE responses with altered conduction properties (as in Sb doping). In all cases, these findings offer a systematic route to achieving spin-control in magnet-free devices, consistent with the broader drive for new chiral spintronics applications [57].
@@ IV Conclusion
In conclusion, this study has demonstrated how substitutional doping in chiral elemental tellurium can systematically tune spin-orbit–driven phenomena such as band splitting, Berry curvature, and the Rashba–Edelstein response. By combining first-principles density functional theory with tight-binding interpolation in PAOFLOW, we revealed that lighter dopants such as sulfur and selenium can shift or reshape the electronic bands near the Fermi level, thereby amplifying collinear spin textures and enhancing components of the Edelstein tensor. In contrast, heavier doping with antimony introduces deeper band restructuring and can either strengthen specific spin-orbit effects or place the Fermi level in different regions of the Brillouin zone, sometimes reducing the net magnitude of current-induced spin polarization. These findings highlight that doping both alters the intrinsic spin-moment locking in chiral crystals and provides a practical means of engineering spin-charge conversion efficiencies without requiring external magnetic fields.
The broader implications of these results extend to magnet-free spintronics, where the ability to selectively enhance or suppress spin relaxation mechanisms is paramount for long-distance spin transport. Because the chirality of tellurium is preserved under moderate doping, the radial or “collinear” spin texture remains robust while undergoing quantitative shifts in spin splitting and curvature hot spots. By choosing dopants of appropriate atomic mass and valence, one can effectively control the band edges and spin splitting, opening a route to optimizing phenomena such as spin Hall effects and radial Edelstein responses. This work thus fills a key gap in the literature, in which the role of doping in chiral semiconductors was previously acknowledged but rarely explored in a systematic ab initio framework.
Beyond the immediate case of tellurium, the methodology employed here is readily transferable to other chiral materials, including the chiral disilicides or selenium-based compounds. Because the computational cost of these simulations remains modest for small supercells and moderate k-point meshes, the approach can be extended to higher dopant concentrations and more complex co-doping scenarios. Such expansions will help further refine our understanding of how spin transport can be tailored through crystal chirality and spin-orbit coupling. Ultimately, these insights can guide the design of next-generation spintronic devices that exploit robust spin accumulation and long spin lifetimes in nonmagnetic systems, thereby reinforcing the importance of chiral semiconductors as a versatile platform for spin-based technologies.

### v3

@@ III.3 Spin texture
The spin textures projected onto the $k_{y}$$k_{z}$ plane reveal marked differences between pristine tellurium (Te) and its element variants (S-, Sb-, and Se-substituted Te). In pristine Te, the spins exhibit a predominantly radial pattern emanating from (or converging toward) high-symmetry points in the Brillouin zone, consistent with the chiral-induced spin textures previously reported for elemental Te. In particular, the spins in pristine Te appear to tilt systematically as one moves along $k_{y}$ or $k_{z}$, a feature that has been attributed to strong spin-orbit coupling (SOC) combined with the crystal’s helical arrangement of atoms. This observation aligns with earlier experimental and theoretical studies of spin textures in chiral Te, which showed similar radial or nearly collinear patterns in momentum space [25, 8]. The net outcome is that, even without an external magnetic field, spin polarization in pristine Te is locked to the crystal momentum in a manner driven by Te’s chirality.
@@
Upon S substitution, the spin texture retains the general radial character but displays a slight enhancement of the spin polarization angles in certain regions of the Fermi surface. In other words, the spins remain roughly outward-pointing, but the degree of out-of-plane tilt varies more strongly across the Brillouin zone. This trend may be attributed to the lighter atomic mass of S relative to Te, subtly modifying the effective SOC strength and shifting the conduction or valence bands. Such element-induced tuning of the spin texture is in line with the view that substituting lighter elements can reduce the overall SOC, thereby sharpening any radial or tangential spin patterns within specific energy pockets [50]. Notably, the S-Te sample maintains the underlying chirality of the host lattice, so the qualitative features of the spin texture remain intact.
@@
By contrast, the Sb-Te system shows a somewhat more pronounced spin polarization, especially at larger $|k_{y}|$ or $|k_{z}|$. This observation suggests that heavier isoelectronic element can amplify SOC-driven spin splitting, leading to larger momentum-dependent variations in spin orientation. The increased atomic mass of Sb relative to Te has often been linked to enhanced spin-orbit coupling, which can in turn reinforce the chiral spin-moment locking in the conduction bands [51]. In the present calculations, the Sb-induced band modifications appear primarily near the Fermi level, thus making the spin texture more conspicuously tilted in those momentum-space regions that dominate conduction. The net result is a element substitution strategy that can, in principle, boost the spin signals relevant for spintronics applications, in keeping with prior theoretical predictions on chiral semiconductors doped with heavier elements [52].
@@
The Se-Substituted Te sample, finally, shows a spin texture that is intermediate between the S and Sb cases, likely reflecting the atomic mass and electronic configuration of Se. Compared to S substituted Te, the Te2Se compound exhibits slightly stronger SOC, manifested as increased tilting in certain $(k_{y},k_{z})$ sectors. However, it lacks the more dramatic spin-splitting enhancements seen in Sb. Such subtle shifts in spin angles and magnitudes may prove critical if one aims to engineer specific transport properties—such as persistent spin helices or large charge-to-spin conversion efficiencies—in a element-dependent fashion [53]. The combination of these four systems underscores how different elements systematically tune the interplay between chirality and SOC, allowing one to either reinforce or diminish the radial-type spin texture.
@@
Taken together, these results corroborate the emerging consensus in the literature that modest element substitution—whether through lighter or heavier substitution—provides a robust means of controlling spin textures in chiral semiconductors [25, 50]. They also reinforce the notion that chiral crystals such as Te enable tunable spin signals even without magnetism, potentially facilitating magnet-free spintronic devices. Consistent with prior first-principles and experimental reports, the radial or nearly collinear spin textures persist across a broad range of elements, demonstrating that the underlying chirality is remarkably resilient. However, the magnitude and tilt of spin polarization can be selectively modified, which paves the way for designing materials with customized spin-transport characteristics. Future work will extend these analyses to higher element concentrations or to co-element scenarios, further mapping out the interplay between element, band-edge positioning, and spin relaxation times in these technologically promising chiral semiconductors.
@@ III.4 The Rashba–Edelstein (REE)
The Rashba–Edelstein (REE) tensor components for pristine Te and three element variants (Te:S, Te:Se, Te:Sb) exhibit clear modifications in both magnitude and energy dependence, underscoring how substitutional doping can tune spin-orbit–driven effects in chiral semiconductors. In pristine Te, the three independent components $\chi_{xx}$, $\chi_{yy}$, $\chi_{zz}$ show intermediate magnitudes on the order of $10^{9}$ (in the chosen units) and prominent dispersion across the studied energy window from about $-0.20$ eV up to 00 eV. Notably, $\chi_{xx}$ and $\chi_{yy}$ overlap partially in magnitude but differ in sign over most of the range, whereas $\chi_{zz}$ exhibits a distinct curvature. These features are consistent with the previously reported “radial” or collinear spin textures in elemental Te, in which the interplay between chiral crystal symmetry and spin-orbit coupling (SOC) leads to a sizable Edelstein response over a broad valence-band region [49, 25].
@@
Upon substituting Te with sulfur (Te:S), the overall magnitude of the tensor components increases to around $10^{10}$, indicating that lighter-element substitution can shift and enhance the SOC-driven effects near the Fermi energy. In particular, $\chi_{xx}$ becomes quite large and negative over a significant portion of the valence region, while $\chi_{yy}$ and $\chi_{zz}$ also remain negative but follow somewhat shallower dispersions. This enhancement is consistent with the notion that element substitution can move the chemical potential to regions where band splittings are pronounced, thereby accentuating the collinear spin polarization [36, 54]. At the same time, the lighter S substitution slightly modifies local bonding environments, which may reduce certain symmetry-allowed band degeneracies and hence enlarge the net REE response.
@@
With selenium (Te:Se) likewise shows magnitudes near $10^{10}$ but with a different balance among the $\chi_{xx}$, $\chi_{yy}$, and $\chi_{zz}$ components. Here, $\chi_{yy}$ acquires its largest negative values around $-0.3$ eV, suggesting that Se substitution emphasizes different orbital characters near the Fermi level compared to S. This behavior is in line with earlier studies that highlight how chalcogen substitutions in Te can shift Weyl-like band features or break certain rotational symmetries, thus affecting both spin-splitting and Berry-related phenomena [55]. The capacity to increase or redirect the REE components through chalcogen substitution corroborates the hypothesis that lighter elements can markedly reshape the conduction- and valence-band edges, thereby offering routes to optimize spin accumulation in nonmagnetic chiral semiconductors [56].
@@
In contrast, with Sb (Te:Sb) reduces the overall scale of the REE tensor to around $10^{7}$. Intriguingly, $\chi_{xx}$ is positive and relatively large compared to $\chi_{zz}$, which is mostly negative throughout the probed energy range. These trends imply that heavier-element substitution, although it enhances intrinsic SOC strength, can simultaneously push the Fermi level away from the band crossings most responsible for large spin-orbit splitting, leading to a net reduction in the Edelstein coefficients [57, 58]. In this sense, Sb might be more suitable for achieving moderate spin-charge conversion while preserving good electronic conduction, rather than maximizing the REE. The interplay between element type (electron vs. hole doping) and element mass therefore emerges as a key strategy for engineering chiral spin transport, aligning with the broader literature on element-induced spin accumulation in noncentrosymmetric semiconductors [49].
@@
Overall, these four cases illustrate how modest substitutional elements in chiral Te can selectively tune the sign, amplitude, and energy dependence of the REE tensor, revealing a clear pathway for tailoring spin-based functionalities in chiral semiconductors. The results agree well with prior theoretical and experimental observations that emphasize element substitution as a pivotal parameter for controlling current-induced magnetization and spin-lifetime effects in topological or near-topological band structures [55, 56]. By situating the Fermi level at different spin-split band extrema—and by exploiting the modified SOC landscapes—one can either amplify spin accumulation (as in S or Se doping) or preserve moderate REE responses with altered conduction properties (as in Sb). In all cases, these findings offer a systematic route to achieving spin-control in magnet-free devices, consistent with the broader drive for new chiral spintronics applications [57].
@@ IV Conclusion
In conclusion, this study has demonstrated how substitutional element in chiral elemental tellurium can systematically tune spin-orbit–driven phenomena such as band splitting, Berry curvature, and the Rashba–Edelstein response. By combining first-principles density functional theory with tight-binding interpolation in PAOFLOW, we revealed that lighter elements such as sulfur and selenium can shift or reshape the electronic bands near the Fermi level, thereby amplifying collinear spin textures and enhancing components of the Edelstein tensor. In contrast, heavier elements such as antimony introduces deeper band restructuring and can either strengthen specific spin-orbit effects or place the Fermi level in different regions of the Brillouin zone, sometimes reducing the net magnitude of current-induced spin polarization. These findings highlight that introducing elements both alters the intrinsic spin-moment locking in chiral crystals and provides a practical means of engineering spin-charge conversion efficiencies without requiring external magnetic fields.
The broader implications of these results extend to magnet-free spintronics, where the ability to selectively enhance or suppress spin relaxation mechanisms is paramount for long-distance spin transport. Because the chirality of tellurium is preserved under moderate element substitution, the radial or “collinear” spin texture remains robust while undergoing quantitative shifts in spin splitting and curvature hot spots. By choosing elements of appropriate atomic mass and valence, one can effectively control the band edges and spin splitting, opening a route to optimizing phenomena such as spin Hall effects and radial Edelstein responses. This work thus fills a key gap in the literature, in which the role of introducing elements in chiral semiconductors was previously acknowledged but rarely explored in a systematic ab initio framework.
Beyond the immediate case of tellurium, the methodology employed here is readily transferable to other chiral materials, including the chiral disilicides or selenium-based compounds. Because the computational cost of these simulations remains modest for small crystals and moderate k-point meshes, the approach can be extended to higher element concentrations and more complex scenarios. Such expansions will help further refine our understanding of how spin transport can be tailored through crystal chirality and spin-orbit coupling. Ultimately, these insights can guide the design of next-generation spintronic devices that exploit robust spin accumulation and long spin lifetimes in nonmagnetic systems, thereby reinforcing the importance of chiral semiconductors as a versatile platform for spin-based technologies.